# Phase Selection and Analysis for Multi-frequency Multi-user RIS Systems Employing Subsurfaces in Correlated Ricean and Rayleigh Environments

### v1

@@ V-B2 Correlation
Fig. 7 compares the mean sum rate for the SD and TMSE designs for a range of element spacings at the RIS and BS. Elements located closer together result in higher correlation. The array parameters are the same as for Figs. 4-6, except for the varied element spacing.
@@
For the SD, it is observed that lower element spacing at the RIS increases rate. This verifies the analysis in Sec. IV-D, as (33) monotonically increases with correlation. In contrast, lower element spacing at the BS decreases rate. This also verifies findings in Sec. IV-D, as (23) decreases with correlation. The RIS element spacing is more influential than the BS element spacing to the rate of the SD. As expected, the rate is higher for the scenario where both spacings are $0.1\lambda$ than when both spacings are $0.5\lambda$. This is due to RIS correlation predominantly affecting the typically larger quadratic $\mathbf{f}_{k}^{\dagger}\mathbf{f}_{k}$ term in (33), and BS correlation mostly affecting the typically smaller $\mathbf{h}_{\mathrm{d}}^{\dagger}\mathbf{f}_{k}$ cross product in (23).
The TMSE method is designed for a RIS-BS channel with a strong LoS component, which is assisted by higher correlation and a higher $\eta_{\mathrm{rb}}$. This is evidenced by smaller RIS element spacing increasing the mean sum rate of the TMSE design. However, as $\eta_{\mathrm{rb}}$ also becomes large, there is a substantial performance degradation. While the algorithm is more accurate for this scenario, strongly LoS channels are not ideal to support multiple users, as the MU channels cannot be separated within a single frequency band. Also, unlike the SD, the BS element spacing is more influential to the rate of the TMSE design. The TMSE method relies on channel separability to serve multiple users in one band. Decreasing the BS element spacing limits this in both the direct and RIS paths, significantly decreasing the performance of the TMSE method. The SD performs comparatively better when correlation is high at the BS, due to separability not being required for its operation.
@@ V-B3 Array Dimensions
M N Mean sum rate decrease from $\eta_{\mathrm{rb}}=0$ to $1$ (%)
@@
M N Mean sum rate increase from $\eta_{\mathrm{rb}}\!=\!0$ to $1$ (bits/s/Hz)
@@
VI Conclusion
In this work, we extend the SD to a higher performing low complexity iterative method (ISD) by setting the subsurfaces sequentially from the weakest user to the strongest so that the users with the largest impact on total SNR benefit from the most prior knowledge. We then propose a converged ISD (CISD), where the ISD is repeated until the SNR increases by less than a specified threshold. We compared the performance of the SD, ISD and CISD, and found that the iterative methods outperform the SD. The most benefit from the CISD occurs at high $N$ where there are more elements that can be further improved with increased knowledge. We also derive the exact closed form expression for the mean SNR of the RIS SD where spatially correlated Ricean fading is assumed for all channels, leading to deeper insights into system performance. This is a very general analysis, and thus we show that the expression can be simplified to verify the results in [12] for correlated Rayleigh UE-BS and UE-RIS channels and a LoS RIS-BS channel. We showed that even while having significantly less BW per user, the SD outperformed the much more complex TMSE method for the realistic scenarios when a channel is strongly LoS or users are clustered together.
Appendix A Derivation of Mean SNR
We compute the mean SNR by taking the expectation of each term in (19). The first two terms, $\mathbb{E}[\mathbf{h}_{\mathrm{d}}^{(k)\dagger}\mathbf{h}_{\mathrm{d}}^{(k)}]$ and $\mathbb{E}[\mathbf{h}_{\mathrm{d}}^{(k)}\mathbf{f}_{k}]$, were derived in [19] with their results listed in (20) and (23). There are four remaining terms to derive. Term 1: $\mathbb{E}[\mathbf{h}_{\mathrm{d}}^{(k)\dagger}\mathbf{g}_{k}]$ As the scattered terms are uncorrelated and hence zero mean,
@@ Appendix A Derivation of Mean SNR
×$\qquad\times\sum_{s\neq k}\mathbf{a}_{\mathrm{r},s}^{(k)\dagger}\mathbf{C}(s)\mathbf{a}_{\mathrm{ur},s}^{(k)},$ (91)
where $\mathbf{C}(s)$ is as given in 29 and derived in Appendix B. Term 2: $\mathbb{E}[\mathbf{f}_{k}^{\dagger}\mathbf{g}_{k}]$ Expanding out $\mathbf{f}_{k}$ and $\mathbf{g}_{k}$ gives
$\mathbb{E}[\mathbf{f}_{k}^{\dagger}\mathbf{g}_{k}]=\mathbb{E}\bigg{[}\mathbf{h}_{\mathrm{ur},k}^{(k)\dagger}\mathbf{\Phi}_{k}^{\dagger}\mathbf{H}_{\mathrm{rb},k}^{(k)\dagger}\sum_{s\neq k}\mathbf{H}_{\mathrm{rb},s}^{(k)}\mathbf{\Phi}_{s}\mathbf{h}_{\mathrm{ur},s}^{(k)}\bigg{]}$
@@
×$\times\mathbb{E}\left[|\mathbf{h}_{\mathrm{ur},k,i}^{(k)}|\right]+\mathbf{R}_{\mathrm{ur},k,s,i,j}^{(k)}\mathbb{E}\left[|\mathbf{h}_{\mathrm{ur},k,i}^{(k)}|^{2}\mathrm{e}^{j\angle\mathbf{h}_{\mathrm{ur},k,i}^{(k)}}\right]\!.$ (96)
@@
[…]
@@
×$\times\int_{0}^{2\pi}\!\!\!\!\mathrm{e}^{j\theta}\mathrm{exp}\!\!\left(\!\!2\sqrt{\frac{\kappa_{\mathrm{ur}}^{(k)}\!\!\left(1+\kappa_{\mathrm{ur}}^{(k)}\right)}{\beta_{\mathrm{ur}}^{(k)}}}r\cos(\theta\!-\!a)\!\!\right)d\theta dr\!.$ (98)
@@
×$\times\!\exp\!\left(\!-\frac{(1+\kappa_{\mathrm{ur}}^{(k)})r^{2}}{\beta_{\mathrm{ur}}^{(k)}}\!\right)\!\!I_{1}\!\!\left(\sqrt{\frac{4\kappa_{\mathrm{ur}}^{(k)}(\kappa_{\mathrm{ur}}^{(k)}+1)}{\beta_{\mathrm{ur}}^{(k)}}}\right)\!\!dr\!.$
@@
×$\!\!\times\!{}_{\!1}{\!F}_{\!1}\!\!\left(\!\frac{3}{2},1,\kappa_{\mathrm{ur}}^{(k)}\!\right)\!\!+\!\frac{3}{2}\mathrm{e}^{j\angle\mathbf{a}_{\mathrm{ur},k,i}^{(k)}}\mathbf{R}_{\mathrm{ur},k,s,i,j}^{(k)}\,{}_{\!1}{\!F}_{\!1}\!\!\left(\!\frac{5}{2},\!1,\!\kappa_{\mathrm{ur}}^{(k)}\!\!\right)\!\!\Bigg{]}\!.\!\!\!$ (100)
@@
Finally, $\mathbf{C}(s)$ is defined in (29) and will be derived in Appendix B. Substituting (95), (100), (Appendix A Derivation of Mean SNR) and (29) into (94) gives (30). Term 3: $\mathbb{E}[\mathbf{f}_{k}^{\dagger}\mathbf{f}_{k}]$ Expanding out $\mathbf{f}_{k}$ with $\mathbf{H}_{\mathrm{rb}}^{(k)}$ split into LoS and scattered components (as seen in (4)) and taking the mean value leaves only the squared terms, due to the independence of the cross product terms. Hence,
@@
×$\!\times\!\exp{\!\left(\!\frac{-2\kappa_{\mathrm{ru}}^{(k)}(1\!-\!\mu_{cf})}{1\!-\!|\rho^{(k)}_{kk}|^{2}}\!\right)}\!\!\sum_{m=0}^{\infty}\sum_{n=0}^{m}\cos(n\phi_{f})\frac{\epsilon_{n}|\rho^{(k)}_{kk}|^{2m-n}}{m!(m\!-\!n)!(n!)^{2}}$
×$\!\times\!\left(\!\frac{\kappa_{\mathrm{ru}}^{(k)}(1+|\rho^{(k)}_{kk}|^{2}\!-\!2\mu_{cf}))}{1-|\rho^{(k)}_{kk}|^{2}}\!\!\right)^{\!\!n}\!\!\Gamma^{2}\!\left(\!m\!+\!\frac{3}{2}\!\right)\!{}_{1}F_{1}^{2}\Bigg{(}\!m\!+\!\frac{3}{2},n\!+\!1,$
@@
Term 4: $\mathbb{E}[\mathbf{g}_{k}^{\dagger}\mathbf{g}_{k}]$ Splitting $\mathbb{E}[\mathbf{g}_{k}^{\dagger}\mathbf{g}_{k}]$ into two sums,
@@
Appendix B Derivation of $\mathbf{C}(x)=\mathbb{E}[\mathbf{\Phi}_{x}]$
@@
Appendix C Derivation of $G_{R}=\mathbb{E}[\mathrm{e}^{-j\phi_{i}}\mathrm{e}^{j\phi_{j}}]$
Let $G_{R}=\mathbb{E}\left[\mathrm{e}^{j\angle\mathbf{h}_{\mathrm{ur},s,i}^{(s)}}\mathrm{e}^{-j\angle\mathbf{h}_{\mathrm{ur},s,j}^{(s)}}\right]=\mathbb{E}[\mathrm{e}^{-j\phi_{i}}\mathrm{e}^{j\phi_{j}}]$. Using the joint PDF for two complex RVs given in (24a) of [20],
[…]
@@ Appendix C Derivation of $G_{R}=\mathbb{E}[\mathrm{e}^{-j\phi_{i}}\mathrm{e}^{j\phi_{j}}]$
×$\times\!\!\!\int_{0}^{\infty}\!\!\!\!\int_{0}^{\infty}\!\!\!r_{\!i}r_{\!j}\!\exp\!\!\left(\!\!\frac{-\big{(}1\!+\!\kappa_{\mathrm{ur}}^{(s)2}\big{)}}{1-\rho_{ss}^{(s)}}\!\!\left(r_{i}^{2}\!+\!r_{j}^{2}\right)\!\!\right)\!\!\int_{-\pi}^{\pi}\!\int_{-\pi}^{\pi}\!\!\mathrm{e}^{-j\phi_{i}}\mathrm{e}^{j\phi_{\!j}}$
×$\times\!\mathrm{e}^{\kappa_{0}r_{i}r_{\!j}\!\cos(\!\phi_{i}\!-\phi_{j}\!-\phi\!)}\mathrm{e}^{\zeta\!\sqrt{\kappa}(r_{i}\cos(\!\phi_{i}\!)+r_{j}\cos(\!\phi_{j}\!)\!)}d\phi_{i}d\phi_{j}dr_{i}dr_{j},$ (131)
@@
We can now integrate with respect to both $\phi_{i}$ and $\phi_{j}$. Rewriting $\mathrm{e}^{-j\phi_{i}}$ and $\mathrm{e}^{j\phi_{j}}$ using Euler’s formula allows the double integral over the phases to be split into four smaller double integrals. Rewriting $\mathrm{e}^{\kappa_{0}r_{i}r_{j}\cos{(\phi_{i}-\phi_{j}-\phi)}}$ in the form seen in [24, 8.511.4] and the resulting $\cos(n\phi_{j}-n\phi_{i}-n\phi)$ in the form seen in [24, 1.313.5], removing periodic terms with zero-valued integrals, applying the product to sum formula found from [24, 1.313.5] and finally integrating each of the four double integrals using [24, 3.915.2], we get
@@
×$\times\!\!\!\int_{0}^{\infty}\!\!\!\!\int_{0}^{\infty}\!\!\!r_{\!i}r_{\!j}\!\exp\!\!\left(\!\!\frac{-\big{(}1\!+\!\kappa_{\mathrm{ur}}^{(s)2}\big{)}}{1-\rho_{ss}^{(s)2}}\!\!\left(r_{i}^{2}\!+\!r_{j}^{2}\right)\!\!\right)\!\!\sum_{n=0}^{\infty}\epsilon_{n}I_{n}(\kappa_{0}r_{i}r_{j})$
×$\times\big{(}J_{n+1}(-j\zeta\sqrt{\kappa}r_{i})J_{n+1}(-j\zeta\sqrt{\kappa}r_{j})\mathrm{e}^{jn\phi}$
@@
where $\epsilon_{0}=1$ and $\epsilon_{n}=2$ when $n\geq 1$. To integrate with respect to $r_{i}$ and $r_{j}$, we can rewrite the modified Bessel function $I_{n}(\kappa_{0}r_{i}r_{j})$ as a Bessel function using [24, 8.406.3], expand that function to its series form using [24, 8.402] and integrate each integral using [24, 6.631.1], resulting in (37).

### v2

@@ V-B2 Correlation
Fig. 8 compares the mean sum rate for the SD and TMSE designs for a range of element spacings at the RIS and BS. Elements located closer together result in higher correlation. As these results involve BS element spacings of $d_{b}=0.1\lambda$, the effects of mutual coupling at the BS must be taken into account. Accordingly, for all curves corresponding to $d_{b}=0.1\lambda$, the mutual coupling model in [29] is incorporated. Specifically, the spatial correlation matrices $\mathbf{R}_{\mathrm{b}}$ and $\mathbf{R}_{\mathrm{d}}$ are replaced by their effective counterparts $\mathbf{R}_{\mathrm{b}}^{\mathrm{eff}}=\mathbf{Z}^{\dagger}\mathbf{R}_{\mathrm{b}}\mathbf{Z}$ and $\mathbf{R}_{\mathrm{d}}^{\mathrm{eff}}=\mathbf{Z}^{\dagger}\mathbf{R}_{\mathrm{d}}\mathbf{Z}$, respectively, where $\mathbf{Z}$ is defined in [29, Eq. (9)]. The array parameters are the same as for Figs. 5-7, except for the varied element spacing.
@@
For the SD, it is observed that lower element spacing at the RIS increases rate. This verifies the analysis in Sec. IV-A, as (23) monotonically increases with correlation. In contrast, lower element spacing at the BS decreases rate. This also verifies findings in Sec. IV-A, as (19) decreases with correlation. The RIS element spacing is more influential than the BS element spacing to the rate of the SD. As expected, the rate is higher for the scenario where both spacings are $0.1\lambda$ than when both spacings are $0.5\lambda$, even when mutual coupling is considered at the BS. This is due to RIS correlation predominantly affecting the typically larger quadratic $\mathbf{f}_{k}^{\dagger}\mathbf{f}_{k}$ term in (23), and BS correlation mostly affecting the typically smaller $\mathbf{h}_{\mathrm{d}}^{\dagger}\mathbf{f}_{k}$ cross product in (19).
The TMSE method is designed for a RIS-BS channel with a strong LoS component, which is assisted by higher correlation and a higher $\eta_{\mathrm{rb}}$. This is evidenced by smaller RIS element spacing increasing the mean sum rate of the TMSE design. However, as $\eta_{\mathrm{rb}}$ also becomes large, there is a substantial performance degradation. While the algorithm is more accurate for this scenario, strongly LoS channels are not ideal to support multiple users, as the MU channels cannot be separated. Also, unlike the SD, the BS element spacing is more influential to the rate of the TMSE design. The TMSE method relies on channel separability to serve multiple users in one band. Decreasing the BS element spacing limits this in both the direct and RIS paths, significantly decreasing the performance of the TMSE method. The SD performs comparatively better when correlation is high at the BS, due to separability not being required for its operation.
@@ V-B3 Array Dimensions
$M$ NN Mean sum rate decrease from $\eta_{\mathrm{rb}}=0$ to $1$ (%)
@@
$M$ NN Mean sum rate increase from $\eta_{\mathrm{rb}}\!=\!0$ to $1$ (bits/s/Hz)
@@
VI Conclusions and Future Work
VI-A Conclusion
In this work, we extended the SD to a higher performing low complexity ISD by setting the subsurfaces sequentially from the weakest user to the strongest so that the users with the largest impact on total SNR benefit from the most prior knowledge. We then proposed the CISD, where the ISD is repeated until the SNR increases by less than a specified threshold. We compared the performance of the SD, ISD and CISD, and found that the iterative methods outperform the SD. The largest benefit from the CISD occurs at high $N$ where there are more elements that can be further improved with increased knowledge. We also derive the exact closed form expression for the mean SNR of the RIS SD where spatially correlated Ricean fading is assumed for all channels, leading to deeper insights into system performance. This is a very general analysis, and thus we show that the expression can be simplified to verify the results in [16] for correlated Rayleigh UE-BS and UE-RIS channels and a LoS RIS-BS channel. We showed that even while having significantly less BW per user, the SD outperformed the much more complex TMSE method for the realistic scenarios when a channel is strongly LoS or users are clustered together.
VI-B Future Work
This work opens up several avenues for future investigation. Firstly, this work focuses on an UL system. The proposed methods are also applicable to DL operation, and under the assumption of perfect CSI, the derived analytical expressions remain structurally unchanged. This assumption is reasonable in the UL, as the required effective channel for RIS phase optimization can be directly inferred from received pilot signals. However, this assumption is less valid in the DL. Relaxing this assumption would require explicitly modeling uncertainty in the channels used for RIS phase design, which would modify the analytical evaluation of the mean SNR. Developing such an analysis constitutes an important direction for future work.
Another important direction for future work is the optimization of the number of RIS elements allocated to each UE. This allocation could be performed either independently of, or jointly with, the RIS phase selection methods proposed in this work. The optimization could be based on a range of criteria, such as each UE’s channel strength, traffic demand, throughput maximization, or fairness objectives. Optimally partitioning the RIS elements among users has the potential to yield significant performance gains and is therefore a promising topic for further research.
Appendix A
Derivation of Mean SNR
We compute the mean SNR by taking the expectation of each term in (17). The first two terms, $\mathbb{E}[\mathbf{h}_{\mathrm{d}}^{(k)\dagger}\mathbf{h}_{\mathrm{d}}^{(k)}]$ and $\mathbb{E}[\mathbf{h}_{\mathrm{d}}^{(k)}\mathbf{f}_{k}]$, were derived in [25] with their results listed in (18) and (19). There are four remaining terms to derive.
Term 1: $\mathbb{E}[\mathbf{h}_{\mathrm{d}}^{(k)\dagger}\mathbf{g}_{k}]$
As the scattered terms are uncorrelated and hence zero mean,
@@ VI-B Future Work
$\qquad\times\sum_{s\neq k}\mathbf{a}_{\mathrm{r},s}^{(k)\dagger}\mathbf{C}(s)\mathbf{a}_{\mathrm{ur},s}^{(k)},$ (38)
where $\mathbf{C}(s)$ is as given in 21 and derived in Appendix B.
Term 2: $\mathbb{E}[\mathbf{f}_{k}^{\dagger}\mathbf{g}_{k}]$
Expanding out $\mathbf{f}_{k}$ and $\mathbf{g}_{k}$ gives
$\mathbb{E}[\mathbf{f}_{k}^{\dagger}\mathbf{g}_{k}]=\mathbb{E}\bigg[\mathbf{h}_{\mathrm{ur},k}^{(k)\dagger}\mathbf{\Phi}_{k}^{\dagger}\mathbf{H}_{\mathrm{rb},k}^{(k)\dagger}\sum_{s\neq k}\mathbf{H}_{\mathrm{rb},s}^{(k)}\mathbf{\Phi}_{s}\mathbf{h}_{\mathrm{ur},s}^{(k)}\bigg],\\$
@@
$\times\mathbb{E}\left[|\mathbf{h}_{\mathrm{ur},k,i}^{(k)}|\right]+\mathbf{R}_{\mathrm{ur},k,s,i,j}^{(k)}\mathbb{E}\left[|\mathbf{h}_{\mathrm{ur},k,i}^{(k)}|^{2}\mathrm{e}^{j\angle\mathbf{h}_{\mathrm{ur},k,i}^{(k)}}\right]\!.$ (41)
@@
$\mathbb{E}\left[r^{2}\mathrm{e}^{j\theta}\right]=\int_{0}^{\infty}\!\!\!\!\int_{0}^{2\pi}r^{2}f^{*}(r,\theta)\,d\theta\,dr,$
@@
$\times\int_{0}^{2\pi}\!\!\!\!\mathrm{e}^{j\theta}\mathrm{exp}\!\!\left(\!\!2\sqrt{\frac{\kappa_{\mathrm{ur}}^{(k)}\!\!\left(1+\kappa_{\mathrm{ur}}^{(k)}\right)}{\beta_{\mathrm{ur}}^{(k)}}}r\cos(\theta\!-\!a)\!\!\right)d\theta dr\!.$ (43)
@@
$\times\!\exp\!\left(\!-\frac{(1+\kappa_{\mathrm{ur}}^{(k)})r^{2}}{\beta_{\mathrm{ur}}^{(k)}}\!\right)\!\!I_{1}\!\!\left(\sqrt{\frac{4\kappa_{\mathrm{ur}}^{(k)}(\kappa_{\mathrm{ur}}^{(k)}+1)}{\beta_{\mathrm{ur}}^{(k)}}}\right)\!\!dr\!.$
@@
$\!\!\times\!{}_{\!1}{\!F}_{\!1}\!\!\left(\!\frac{3}{2},1,\kappa_{\mathrm{ur}}^{(k)}\!\right)\!\!+\!\frac{3}{2}\mathrm{e}^{j\angle\mathbf{a}_{\mathrm{ur},k,i}^{(k)}}\mathbf{R}_{\mathrm{ur},k,s,i,j}^{(k)}\,{}_{\!1}{\!F}_{\!1}\!\!\left(\!\frac{5}{2},\!1,\!\kappa_{\mathrm{ur}}^{(k)}\!\!\right)\!\!\Bigg]\!.\!\!\!$ (45)
@@
Finally, $\mathbf{C}(s)$ is defined in (21) and will be derived in Appendix B. Substituting (40), (45), (Appendix A Derivation of Mean SNR) and (21) into (39) gives (22).
Term 3: $\mathbb{E}[\mathbf{f}_{k}^{\dagger}\mathbf{f}_{k}]$
Expanding out $\mathbf{f}_{k}$ with $\mathbf{H}_{\mathrm{rb}}^{(k)}$ split into LoS and scattered components (as seen in (4)) and taking the mean value leaves only the squared terms, due to the independence of the cross product terms. Hence,
@@
$\!\times\!\exp{\!\left(\!\frac{-2\kappa_{\mathrm{ru}}^{(k)}(1\!-\!\mu_{cf})}{1\!-\!|\rho^{(k)}_{kk}|^{2}}\!\right)}\!\!\sum_{m=0}^{\infty}\sum_{n=0}^{m}\cos(n\phi_{f})\frac{\epsilon_{n}|\rho^{(k)}_{kk}|^{2m-n}}{m!(m\!-\!n)!(n!)^{2}}$
$\!\times\!\left(\!\frac{\kappa_{\mathrm{ru}}^{(k)}(1+|\rho^{(k)}_{kk}|^{2}\!-\!2\mu_{cf}))}{1-|\rho^{(k)}_{kk}|^{2}}\!\!\right)^{\!\!n}\!\!\Gamma^{2}\!\left(\!m\!+\!\frac{3}{2}\!\right)\!{}_{1}F_{1}^{2}\Bigg(\!m\!+\!\frac{3}{2},n\!+\!1,$
@@
Term 4: $\mathbb{E}[\mathbf{g}_{k}^{\dagger}\mathbf{g}_{k}]$
Splitting $\mathbb{E}[\mathbf{g}_{k}^{\dagger}\mathbf{g}_{k}]$ into two sums,
@@
Appendix B
Derivation of $\mathbf{C}(x)=\mathbb{E}[\mathbf{\Phi}_{x}]$
@@
Appendix C
Derivation of $G_{R}=\mathbb{E}[\mathrm{e}^{-j\phi_{i}}\mathrm{e}^{j\phi_{j}}]$
Let $G_{R}=\mathbb{E}\left[\mathrm{e}^{j\angle\mathbf{h}_{\mathrm{ur},s,i}^{(s)}}\mathrm{e}^{-j\angle\mathbf{h}_{\mathrm{ur},s,j}^{(s)}}\right]=\mathbb{E}[\mathrm{e}^{-j\phi_{i}}\mathrm{e}^{j\phi_{j}}]$. Using the joint PDF for two complex random variables (RVs) given in (24a) of [26],
$G_{R}\!=\!\!\!\int_{0}^{\infty}\!\!\!\!\int_{0}^{\infty}\!\!\!\!\int_{-\pi}^{\pi}\!\int_{-\pi}^{\pi}\!\!\!\!\mathrm{e}^{\!-j\phi_{i}}\mathrm{e}^{j\phi_{j}}\!f^{*}\!\!\left(r_{\!i},r_{\!j},\mathrm{e}^{j\phi_{i}}\!,\mathrm{e}^{j\phi_{j}}\!\right)\!d{\phi_{i}}d{\phi_{j}}d{r_{\!i}}d{r_{\!j}}.$
@@
$\times\!\!\!\int_{0}^{\infty}\!\!\!\!\int_{0}^{\infty}\!\!\!r_{\!i}r_{\!j}\!\exp\!\!\left(\!\!\frac{-\big(1\!+\!\kappa_{\mathrm{ur}}^{(s)2}\big)}{1-\rho_{ss}^{(s)}}\!\!\left(r_{i}^{2}\!+\!r_{j}^{2}\right)\!\!\right)\!\!\int_{-\pi}^{\pi}\!\int_{-\pi}^{\pi}\!\!\mathrm{e}^{-j\phi_{i}}\mathrm{e}^{j\phi_{\!j}}$
$\times\!\mathrm{e}^{\kappa_{0}r_{i}r_{\!j}\!\cos(\!\phi_{i}\!-\phi_{j}\!-\phi\!)}\mathrm{e}^{\zeta\!\sqrt{\kappa}(r_{i}\cos(\!\phi_{i}\!)+r_{j}\cos(\!\phi_{j}\!)\!)}d\phi_{i}d\phi_{j}dr_{i}dr_{j},$ (66)
@@
We can now integrate with respect to both $\phi_{i}$ and $\phi_{j}$. Rewriting $\mathrm{e}^{-j\phi_{i}}$ and $\mathrm{e}^{j\phi_{j}}$ using Euler’s formula allows the double integral over the phases to be split into four smaller double integrals. Rewriting $\mathrm{e}^{\kappa_{0}r_{i}r_{j}\cos{(\phi_{i}-\phi_{j}-\phi)}}$ in the form seen in [31, 8.511.4] and the resulting $\cos(n\phi_{j}-n\phi_{i}-n\phi)$ in the form seen in [31, 1.313.5], removing periodic terms with zero-valued integrals, applying the product to sum formula found from [31, 1.313.5] and finally integrating each of the four double integrals using [31, 3.915.2], we get
@@
$\times\!\!\!\int_{0}^{\infty}\!\!\!\!\int_{0}^{\infty}\!\!\!r_{\!i}r_{\!j}\!\exp\!\!\left(\!\!\frac{-\big(1\!+\!\kappa_{\mathrm{ur}}^{(s)2}\big)}{1-\rho_{ss}^{(s)2}}\!\!\left(r_{i}^{2}\!+\!r_{j}^{2}\right)\!\!\right)\!\!\sum_{n=0}^{\infty}\epsilon_{n}I_{n}(\kappa_{0}r_{i}r_{j})$
$\times\big(J_{n+1}(-j\zeta\sqrt{\kappa}r_{i})J_{n+1}(-j\zeta\sqrt{\kappa}r_{j})\mathrm{e}^{jn\phi}$
@@
where $\epsilon_{0}=1$ and $\epsilon_{n}=2$ when $n\geq 1$. To integrate with respect to $r_{i}$ and $r_{j}$, we can rewrite the modified Bessel function $I_{n}(\kappa_{0}r_{i}r_{j})$ as a Bessel function using [31, 8.406.3], expand that function to its series form using [31, 8.402] and integrate each integral using [31, 6.631.1], resulting in (27).